# Opening Articulated Objects in the Real World

### v2

@@ VI Discussion
• Perception remains the biggest bottleneck in deploying such a mobile manipulation system to in-the-wild settings. Perception failures (Mask RCNN, Detic) reveal the need for broader datasets spanning diverse environments (i.e. everyday cluttered offices / apartments beyond luxury homes). This highlights the importance of prioritizing advancements in perception modules to accelerate progress toward generalizable mobile manipulation systems.
• While the system demonstrated robustness to inaccuracies in articulation parameter estimation, successfully opening cabinets even with a 10cm error in radius (Section IV-D), achieving an initial secure grasp remained a critical challenge. Grasping failures accounted for 25% of all failures, highlighting the difficulty of precise last-centimeter grasping. Contact correction provided a partial remedy, but closed-loop visual grasping strategies of some form that can compensate for minor mis-calibrations and imprecise navigation could help further.
@@
A7 Supporting Figures and Tables
@@
A8 Evaluation of Motion Plan Generation
@@
A9 Detic [70]-based Articulation-parameter Prediction Module
@@
A10 Other Articulation Types
@@
A11 Implementation Details for Baselines
@@
A11-A Robot Utility Models
@@
A11-B Sim2Real Behavior Cloning
@@
A11-C Real World Behavior Cloning
@@
A11-E Other Prior Systems
@@
A12 Comparisons to Prior Work on Articulation Parameter Estimation
@@
A13 Comparison to Direct 3D Prediction

### v3

@@ VI Discussion
One might wonder how we contextualize our paper with respect to the Bitter Lesson. The Bitter Lesson argues that general-purpose methods that scale with data and compute (e.g. end-to-end learning) ultimately outperform systems built with handcrafted structure. However, our work provides concrete counter evidence: MOSART is a modular system that outperforms RUM [16], a state-of-the-art imitation learning (IL) system by an absolute 35%. Note, this IL system was already scaled-up: it used 1000+ demos for training. With enough data end-to-end learning may prevail, but unlike language and vision domains where data is plentiful, robotics faces unique challenges that raise questions about whether we will ever achieve the data volume needed for true generalization. Simulation could help, but the sim-to-real gap remains a major obstacle.
Furthermore, the bitter lesson may not always hold. For example, the success of NeRFs [50] entirely relies on embedding volumetric rendering into learning—perhaps, modularity is a similar structure to make robotic systems effective. There is increasing evidence pointing in this direction: [22] shows that a modular system outperforms a scaled-up IL/RL system for navigation, while [29] separates perception + high-level reasoning from low-level whole-body control for humanoid control. Furthermore, a modular system doesn’t have to be anti-scaling. Rather, modularity can be very much pro-scaling: each individual module can be scaled up to absorb much more data than a monolithic end-to-end system. For example, the perception module in MOSART is trained on Internet images, far surpassing the data and diversity that an IL system would be exposed to. These observations suggest that modularity may not only be compatible with scale, but could be essential for achieving generalization in real world robotics.
@@
• Perception remains the biggest bottleneck in deploying such a mobile manipulation system to in-the-wild settings. The perception failures, even with strong models such as Mask RCNN and Detic, reveal a critical need for broader and more diverse datasets that better capture the variability of real world environments (i.e. everyday cluttered offices / apartments beyond luxury homes). This highlights the importance of prioritizing advancements in perception modules to accelerate progress toward generalizable mobile manipulation systems.
• While the system demonstrated robustness to inaccuracies in articulation parameter estimation, successfully opening cabinets even with a 10cm error in radius (Section IV-D), achieving an initial secure grasp of the handle remained a critical challenge. Grasping failures accounted for approximately 25% of all observed failures, underscoring the inherent difficulty of achieving precise, last-centimeter adjustments required for successful grasping. Contact correction provided a partial remedy, but closed-loop visual grasping strategies of some form that can compensate for minor mis-calibrations and imprecise navigation could help further.
@@
VII Acknowledgments
This material is based upon work supported by DARPA (Machine Common Sense program), an NSF CAREER Award (IIS-2143873), and the Andrew T. Yang Research and Entrepreneurship Award. We are grateful to the Centre for Autonomy for lending us the Stretch RE2 robot used in this work. We thank Aditya Prakash and Matthew Chang for their feedback on manuscript.
@@ VII Acknowledgments
A8 Supporting Figures and Tables
@@
A9 Evaluation of Motion Plan Generation
@@
A10 Detic [89]-based Articulation-parameter Prediction Module
@@
A11 Other Articulation Types
@@
A12 Implementation Details for Baselines
@@
A12-A Robot Utility Models
@@
A12-B Sim2Real Behavior Cloning
@@
A12-C Real World Behavior Cloning
@@
A12-E Other Prior Systems
@@
A13 Timing Comparison to RUM
MOSART is significantly faster than RUM [16]. It executes in 80.68 sec vs. 129.68 sec taken by RUM. MOSART ’s 80.68 secs break down as: perception (7.61 sec), transforms (0.319 sec), navigation (17.33 sec), motion planning (0.091 sec), executing the pre-grasp pose (19 sec), contact correction (12 sec), and the full execution (24.33 sec). Compute is done locally on the Stretch’s Intel i5 processor.
A14 Comparisons to Prior Work on Articulation Parameter Estimation
@@
A15 Comparison to Direct 3D Prediction
@@
A16 Comparison to Grounded SAM 2
Additionally, we also compare to Grounded SAM 2 [67], and observe similar trends as Detic (see Table II): slightly better mean handle error but higher mean radius error due to occasional clumping of multiple cabinets (e.g. Figure A11). Detection, orientation, and articulation accuracies remain high.
Grounded SAM 2 Ours
Detection 29/31 29/31
Handle orientation 29/29 28/29
Articulation type 28/29 29/29
Mean handle error 1.34 cm 2.11 cm
Mean radius error 7.71 cm 0.98 cm
Table A6: Comparison to Grounded SAM 2.